A plane can fly 480 miles in the same time as it takes a car to go 210 miles. If the car travels 90 mph slower than the plane, find the speed of the plane.
Answer by scott8148(6628)

t = d / r

480 / p = 210 / (p - 90)

"cross" multiplying ___ 480p - (90 * 480) = 210p

270p = 90 * 480